OpenStudy (anonymous):

Evaluate the double integral S(-5)--->(6)S(10)--->12 (1/(x^2-3x+2))dxdy using partial fractions

myininaya (myininaya):

\[\int\limits_{}^{}\int\limits_{}^{}\frac{1}{x^2-3x+2} dx dy \] I don't understand S(-5)--->(6)S(10)--->12 I have no clue what that means.

myininaya (myininaya):

I can help you on the partial fraction part though.

myininaya (myininaya):

You should be able to factored x^2-3x+2

OpenStudy (anonymous):

That is what the integrals go from. The first integrals range (dx) is 10 to 12 and the second integrals range is 5 to 6.

myininaya (myininaya):

So you mean for those S's to be integral signs.

myininaya (myininaya):

But anyways let's factor x^2-3x+2 That is the first step

OpenStudy (anonymous):

When i did the partial fractions, I got -2(x-1)^-1 + 2(x-2)^-1

OpenStudy (anonymous):

Yes they were supposed to be integral signs. haha

myininaya (myininaya):

\[\frac{1}{x^2-3x+2}=\frac{A}{(x-2)}+\frac{B}{(x-1)}\] Combine those fractions we get: \[\frac{1}{x^2-3x+2}=\frac{A(x-1)+B(x-2)}{(x-2)(x-1)}\] Since we have the bottoms the sign and we want this to be the same fraction then we need to find for what A and B does 1=A(x-1)+B(x-2) 1=(A+B)x+(-A-2B) There are no x's on that one side so we know A+B=0 We have the constant 1 on that one side so we know we need -A-2B=1 So is this what you did to get A and B?

myininaya (myininaya):

I'm not getting those 2's you got.

myininaya (myininaya):

Do you see what I mean?

OpenStudy (anonymous):

Thats how I did it, but when I solved for B i got 2.

OpenStudy (anonymous):

THe other two was a mistake.

myininaya (myininaya):

So A+B=0 => -A=B right?

myininaya (myininaya):

In our second equation we have -A-2B=1

myininaya (myininaya):

Replace -A with B.

myininaya (myininaya):

Solve for B.

OpenStudy (anonymous):

B-2B=-1

myininaya (myininaya):

Well B-2B=1

myininaya (myininaya):

So -1B=1 which implies B=-1

myininaya (myininaya):

A should be easy to find now since you know that A and B are opposites of each other.

OpenStudy (anonymous):

A=1

myininaya (myininaya):

Ok so we have this now: \[\int\limits_{-5}^{6}\int\limits_{10}^{12}( \frac{1}{x-2}+\frac{-1}{x-1}) dx dy\]

myininaya (myininaya):

You can find what I'm about to write by substitution or you can remember this formula for the antiderivative for 1/(x-c). \[\int\limits_{}^{}\frac{1}{x-c}dx=\ln|x-c|+k \text{ where c is a constant } \]

OpenStudy (anonymous):

ok so after the first integral i got (ln 11 - ln 10)-(ln9-ln8)

myininaya (myininaya):

uhhh. do i have the intervals backwards or do you ?

OpenStudy (anonymous):

no its correct.

myininaya (myininaya):

So 10 to 12 not 12 to 10 right?

myininaya (myininaya):

\[[\ln|x-2|-\ln|x-1|]_{10}^{12}=(\ln(10)-\ln(11))-(\ln(8)-\ln(9))\]

OpenStudy (anonymous):

ya, andthen after the second integral it would just be equal to that right?

myininaya (myininaya):

Well no...

myininaya (myininaya):

\[\int\limits_{-5}^{6}c dy=cy|_{-5}^{6}=c(6)-c(-5)=6c+5c\]

myininaya (myininaya):

or 11c

OpenStudy (anonymous):

oops i meant 5 to 6. Not -5 to 6...